OpenStudy (anonymous):

Factor completely 64x^6 - y^6

OpenStudy (anonymous):

64x^6 = ((2x)^3)^2 y^6 = (y^3)^2 take them as a and b . you look up the formula at http://www.engineeringtoolbox.com/factors-expansion-d_1249.html and you can have it done by yourself. I give you the way to do only. I don't do your homework. sorry for that

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Thanks